OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (anthonyym):

What is the question?

OpenStudy (abbycross167):

Please help me to simplify this rational expression - (6x) (2y+2) ------ * ------- (y+1) (3)

jhonyy9 (jhonyy9):

(3*2x) 2(x+1) ----- * -------- (x+1) 3 so do you see it now what and how can simplifie there ?

OpenStudy (anthonyym):

@jhonyy9 I don't understand what you did because the y's disappeared

OpenStudy (abbycross167):

I'm confused as well... @jhonyy9 could you try to explain it how to do it? Please?

jhonyy9 (jhonyy9):

(6x) (2y+2) ----- * ------- so you can facvtorizing out and rewrite the 6 in this way (y+1) 3 3*2x 2(y+1) ----- * ------- = (y+1) 3 hopw you see now clearly what can you easy cross simplifie there

jhonyy9 (jhonyy9):

factorizing out the 2 from there

jhonyy9 (jhonyy9):

do you see it now what can you cross simplifie easy ?

jhonyy9 (jhonyy9):

please collaborate from 3*2x the 3 and cros to this the 3 too - yes ?

OpenStudy (anthonyym):

Oh I see now. You mean 6x 2(y+1) ----* ----- y+1 3

jhonyy9 (jhonyy9):

yes so 3 from 3*2x with 3 from denominator simplifie it sure easy and from the laft denominator the (y+1) with (y+1) from numerator on the right side

OpenStudy (anthonyym):

@abbycross167 We factored a 2 out from 2y+2 so that it becomes 2(y+1). So we have \[\frac{ 6x }{ y+1 }*\frac{ 2(y+1) }{ 3 }\] Now we can cancel out the y+1.

OpenStudy (abbycross167):

@jhonyy9 @anthoyym I am so sorry I haven't been on here to help!!! My computer shut off and will not turn back on, I'm very sorry...

jhonyy9 (jhonyy9):

np good luck

OpenStudy (abbycross167):

@jhonyy9 would you mind helping me still?